The length of a rectangle is 2x and its area is 6x^2. Find its perimeter.

Please tell me the steps on how to solve it. Thanks

Answer by lwsshak3(11628)

The length of a rectangle is 2x and its area is 6x^2. Find its perimeter
**
length times width=area
2x times width=6x^2
width=Area/length=6x^2/2x=3x
Dimensions of rectangle is 2x by 3x
Perimeter=2(length+width)=2(2x+3x)=10x